# Predictively Oriented Posteriors

### v1

@@ Appendix D Further experimental details
D.1 Regression with the MMD
sec:synthetic-regression-supp Denote $d=\operatorname{dim}(\Theta)$. In Section 5.2 we use the squared MMD loss to fit the linear regressions, $\{\mathsf{N}(\cdot;\,x^{\top}\theta,\sigma^{2}):\,x\in\mathbb{R}^{d},\,\theta\in\mathbb{R}^{d}\}$, where $\sigma^{2}>0$ is treated as fixed and is estimated from data:
$\hat{\sigma}^{2}=\sum_{i=1}^{n}\frac{(y_{i}-\hat{y}_{i})}{n-d-1};\quad\hat{y}=X\hat{\beta};\quad\hat{\beta}=(X^{\top}X)^{-1}X^{\top}y.$
In the case of the Gibbs posterior, we use the efficient estimator proposed by Alquier and Gerber (2024, Equation 6):
$S_{n}(P_{\theta})=\sum_{i=1}^{n}\ell^{\dagger}(\theta;\,X_{i},y_{i})$
where for all $\theta\in\Theta,\,x\in\mathbb{R}^{d},$ and $y\in\mathbb{R}$,
$\ell^{\dagger}(\theta;\,X,y)=\mathbb{E}_{Y,Y^{\prime}\sim P_{\theta}(\cdot\mid x)}\{\kappa(Y,Y^{\prime})-2\kappa(Y,y)\}.$
In practice we estimate both of these expectations with $m$ Monte Carlo draws:
$\ell^{\dagger}(\theta;\,X,y)\approx\frac{1}{m(m-1)}\sum_{i\neq j}\sum_{j=1}^{m}\kappa(Y_{i},Y^{\prime}_{j})-\frac{2}{m}\kappa(Y_{i},y),$
for $Y_{i},Y^{\prime}_{i}\sim P_{\theta}(\cdot\mid x).$
In the case of the PrO-posterior, we instead use the tandem loss
$L_{n}(\theta,\theta^{\prime})=\sum_{i=1}^{n}\ell(\theta,\theta^{\prime};\,X_{i},y_{i})$
where now
$\ell(\theta,\theta^{\prime};\,X,y)=\mathbb{E}_{\begin{subarray}{c}Y\sim P_{\theta}(\cdot\mid x)\\ Y^{\prime}\sim P_{\theta^{\prime}}(\cdot\mid x)\end{subarray}}\{\kappa(Y,Y^{\prime})-\kappa(Y,y)-\kappa(Y^{\prime},y)\}.$
And again, each of these expectations can be approximated by Monte Carlo:
$\ell(\theta,\theta^{\prime};\,X,y)\approx\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}\kappa(Y_{i},Y^{\prime}_{j})-\frac{1}{m}\kappa(Y_{i},y)-\frac{1}{m}\kappa(Y_{i}^{\prime},y),$
for $Y_{i}\sim P_{\theta}(\cdot\mid x)$ and $Y^{\prime}_{i}\sim P_{\theta^{\prime}}(\cdot\mid x)$.
In both cases, we use a Gaussian kernel where the length scale is chosen according to the median heuristic (Garreau et al., 2018).
@@ D.2 Palmer penguins example
The data containing $n=342$ measurements of the penguins’ bill lengths and depths in millimetres were imported from the palmerpenguins package in Python, and jointly centred to have zero-mean and scaled to have unit variance. To these data we fit a Gaussian kernel density estimator implemented in sklearn (Pedregosa et al., 2011), where the bandwidth was chosen with cross-validation across a logarithmic grid. We also fit for comparison a PrO posterior over the mean of an the isotropic bivariate Gaussian location model class, $\mathcal{M}_{\Theta}=\{\mathsf{N}(\theta,\sigma^{2}I):\,\theta\in\mathbb{R}^{2}\}$, where $\sigma^{2}$ is fixed, and which we fit with the squared MMD score. This same bandwidth was used for fitting the PrO posterior with the squared MMD as fitting the KDE. We estimated the squared MMD with Monte Carlo using $m=n$ samples (in the notation of Section D.1), $\lambda_{n}=10^{3}$, and $\sigma^{2}=0.2$. The Wasserstein gradient flow was run with $2^{5}$ particles and $\mathsf{d}t=10^{-1}$ for $10,000$ iterations. In this example, and in all following examples, both Gibbs and PrO posteriors are fit with the same value of $\lambda_{n}$.
D.3 Normal location illustrations
We simulate $n=1,000$ data from each of the data-generating processes:
                                                                                                    $y_{i}$ ${}\sim{}$ $\mathsf{N}(0,1^{2})$ (“Well-specified”)
                                                                                                    $y_{i}$ ${}\sim{}$ $\begin{cases}\mathsf{N}(-2,1^{2}),&\quad\text{with probability }0.2\\ \mathsf{N}(2,1^{2}),&\quad\text{with probability }0.8\end{cases}$ (“Mixture”)
@@ D.3 Normal location illustrations
                                                                                                    $y_{i}$ ${}\sim{}$ $\mathsf{Student}-t_{1.5}$ $\quad\text{(``Heavy-tails'')}.$
For each of these, we estimate the MMD with $m=2^{7}$ Monte Carlo samples, run the Wasserstein gradient flow for $p=2^{5}$ particles for all data, and for $4,000$ iterations similarly to MALA. We use $\lambda_{n}=n$. The values of $\mathsf{d}t$ for the WGF and MALA varied between datasets and were chosen by visual inspection of the trajectory plots.
Figure 8: Illustrative examples of Section 1.2. In the left-most column we show the particle trajectories from a Wasserstein gradient flow targeting the PrO-posterior (in blue) compared to the path of a single MALA chain targeting the Gibbs posterior (in red). In the two axes on the right we show the empirical posterior densities from these two trajectories.
As mentioned before, the predictive distribution in the case of the normal location model is nothing more than a smoothed posterior distribution. We show the posterior distributions of the Gibbs and PrO-posteriors and the Wasserstein gradient flow trajectories in Figure 8.
D.4 Golf putting example
We set $\lambda_{n}=n^{1/2}$ and approximate the log score with a DI approximation as described by Masegosa (2020). The WGF was run with different step sizes for both dimensions, $\mathsf{d}t=(8\times 10^{-4},2\times 10^{-5})$, and with $p=10$ particles for $8,000$ iterations. MALA was run with $\mathsf{d}t=3\times 10^{-7}$ and also for $8,000$ iterations. The initial particle values for both methods were initialised around the posterior mean values (2.23,0.26)(2.23,-0.26) identified by Gelman and Nolan (2002).
@@ D.5 Binary classification example
We simulated $n=1,000$ in the manner described in Section 5.1, and fit all models with $\lambda_{n}=n^{1/2}$. The PrO posterior was based on the log score, and fit under two different approximations: the same DI approximation as in Section 5.3, and an MS approximation with $k=2$. Both were run with $p=2^{6}$ particles, $\mathsf{d}t=5\times 10^{-2}$, and for $10,000$ iterations. MALA was run for $4,000$ steps with $\mathsf{d}t=10^{-4}$.
D.6 Synthetic linear regression
In Section 5.2 we set $\xi=2,\,\sigma^{2}=1,$ and $\eta=0.5$ to generate $n=500$ training data. The log point-wise predictive density (lppd) is estimated from $S$ samples from the posterior $\{\theta^{(s)}\}_{s=1}^{S}$ at a point $\{X_{i},y_{i}\}$ through
$\operatorname{lppd}_{i}=\log\left\{\frac{1}{S}\sum_{s=1}^{S}P_{\theta^{(s)}}(y_{i}\mid X_{i})\right\}.$
In the right-most panel of Figure 6 we plot the difference $\Delta_{i}=\operatorname{lppd}_{i}^{(\operatorname{PrO})}-\operatorname{lppd}_{i}^{(\operatorname{Gibbs})}$. And since we want to maximise the log score, if $\Delta_{i}>0$ we say that the PrO posterior induces better predictive performance. The expected log predictive density (elpd) on a test set, $\{X_{i},y_{i}\}_{i=1}^{n_{\mathrm{test}}}$ is computed as
$\operatorname{elpd}=\sum_{i=1}^{n_{\mathrm{test}}}\operatorname{lppd}_{i}.$
The PrO posterior attains an elpd of 1834.68-1834.68 on $n_{\mathsf{test}}=1,000$ test data, while the Gibbs posterior performs much worse with an elpd of 3593.67-3593.67 on the same data.
The MMD was estimated using $m=2^{5}$ Monte Carlo samples, and the Wasserstein gradient flow was run with $\mathsf{d}t=5\times 10^{-3}$ using $p=2^{4}$ particles and $\lambda_{n}=n^{1/2}$ for $4,000$ iterations, while MALA was run with $\mathsf{d}t=10^{-4}$ also for $4,000$ iterations after $96,000$ warmup iterations to ensure convergence.

### v2

@@ Appendix D Further experimental details
D.1 Normal location illustrations
[…]
@@ D.2 Palmer penguins example
The data containing $n=342$ measurements of the penguins’ bill lengths and depths in millimetres were imported from the palmerpenguins package in Python, and jointly centred to have zero-mean and scaled to have unit variance. To these data we fit a Gaussian kernel density estimator implemented in sklearn (Pedregosa et al., 2011), where the bandwidth was chosen with cross-validation across a logarithmic grid. We also fit for comparison a PrO posterior over the mean of an the isotropic bivariate Gaussian location model class, $\mathcal{M}_{\Theta}=\{\mathsf{N}(\theta,\sigma^{2}I):\,\theta\in\mathbb{R}^{2}\}$, where $\sigma^{2}=0.2$ is fixed, and which we fit with the squared MMD score. This same bandwidth was used for fitting the PrO posterior with the squared MMD as fitting the KDE. We estimated the PrO posterior under the squared MMD loss approximated with $m=n$ samples, $\lambda_{n}=10^{3}$, and the Wasserstein gradient flow was run with $p=32$ particles for $10,000$ iterations.
D.3 Regression with the MMD
Denote $d=\operatorname{dim}(\Theta)$. In Sections 25.2, and 5.3 we fit the linear models under the squared MMD loss. Denote the model class as $\{\mathsf{N}(y_{i};\,x_{i}^{\top}\theta,\sigma^{2}):\,\theta\in\mathbb{R}^{d}\}$, for $i=1,\ldots,n$ where $\sigma^{2}>0$ is treated as fixed and estimated from data: $\hat{\sigma}^{2}=\sum_{i=1}^{n}\frac{(y_{i}-\hat{y}_{i})}{n-d-1};\quad\hat{y}=X\hat{\beta};\quad\hat{\beta}=(X^{\top}X)^{-1}X^{\top}y.$ In the case of the Gibbs posterior, we use the efficient estimator proposed by Alquier and Gerber (2024, Equation 6): $S_{n}(P_{\theta})=\sum_{i=1}^{n}\ell^{\dagger}(\theta;\,X_{i},y_{i})$ where for all $\theta\in\Theta,\,x\in\mathbb{R}^{d},$ and $y\in\mathbb{R}$, $\ell^{\dagger}(\theta;\,X,y)=\mathbb{E}_{Y,Y^{\prime}\sim P_{\theta}(\cdot\mid x)}\{\kappa(Y,Y^{\prime})-2\kappa(Y,y)\}.$ In practice we estimate both of these expectations with $m$ Monte Carlo draws: $\ell^{\dagger}(\theta;\,X,y)\approx\frac{1}{m(m-1)}\sum_{i\neq j}\sum_{j=1}^{m}\kappa(Y_{i},Y^{\prime}_{j})-\frac{2}{m}\kappa(Y_{i},y),$ for $Y_{i},Y^{\prime}_{i}\sim P_{\theta}(\cdot\mid x).$ In the case of the PrO posterior, we instead use the tandem loss $L_{n}(\vartheta,\theta)=\sum_{i=1}^{n}\ell(\vartheta,\theta;\,X_{i},y_{i})$ where now $\ell(\vartheta,\theta;\,X,y)=\mathbb{E}_{\begin{subarray}{c}Y\sim P_{\vartheta}(\cdot\mid x)\\ Y^{\prime}\sim P_{\theta}(\cdot\mid x)\end{subarray}}\{\kappa(Y,Y^{\prime})-\kappa(Y,y)-\kappa(Y^{\prime},y)\}.$ And again, each of these expectations can be approximated by Monte Carlo: $\ell(\vartheta,\theta;\,X,y)\approx\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}\kappa(Y_{i},Y^{\prime}_{j})-\frac{1}{m}\kappa(Y_{i},y)-\frac{1}{m}\kappa(Y_{i}^{\prime},y),$ for $Y_{i}\sim P_{\vartheta}(\cdot\mid x)$ and $Y^{\prime}_{i}\sim P_{\theta}(\cdot\mid x)$. In both cases, we use a Gaussian kernel where the length scale is chosen according to the median heuristic (Garreau et al., 2018).
D.4 Linear regression example
@@
D.4.1 Sensitivity of the Wasserstein gradient flow
In Figure 14 we show the sensitivity of the kernel gradient discrepancy (a diagnostic of the quality of the posterior samples) and two predictive scores to the choice of the number of particles and step size in a linear regression example. We sample $n=500$ data from the CR setting described in Section 2, and then fit the PrO posterior with $p=\{2^{1},2^{2},\ldots,2^{6}\}$ particles and $\mathsf{d}t$ between $0.1$ and $0.0001$. In each case, we set $\lambda_{n}=n^{1/2}$ and run the WGF for $10,000$ iterations and estimate the MMD with $m=32$ Monte Carlo samples. In Figure 14 we report the mean of the kernel gradient discrepancy (KGD), and predictive negative log likelihood (NLL) and MMD from $10$ fold cross-validation. All metrics are improved as $\mathsf{d}t$ is smaller, and as the number of particles grows. We observed here, and in other experiments, that as long as the number of particles is sufficiently large to capture multi-modal behaviour in the data, the step size becomes important to tune appropriately. And that the relationship relationship between predictive performance and step size is not always linear. This observation led us to choose $p$ of the order $\approx 2^{5}$, and optimise the step size in all experiments with the tuning-free FUSE schedule (Sharrock and Nemeth, 2025).
Refer to caption (a) Posterior diagnostic.
Refer to caption (b) Predictive metrics.
Figure 14: Synthetic linear regression example.
@@ D.5 Binary classification example
We simulated $n=1,000$ in the manner described in Section 5.1, and fit all models with $\lambda_{n}=n^{1/2}$. The PrO posterior was run with $p=64$ particles for $10,000$ iterations. MALA was run for $4,000$ steps with $\mathsf{d}t=10^{-4}$.
D.6 River water flow example
We fit an $\operatorname{AR}(7)$ model with Student-$t$ prior over the intercept $\beta$, and Minnesota-style priors over the regression coefficients $\theta$:                                                                                                     $\Pi(\beta)$ ${}\sim{}$ $t_{3}(0,1^{2})$                                                                                                                                                                                                          $\Pi(\theta_{j})$ ${}\sim{}$ $\mathsf{N}\left(0,\,\frac{1}{j^{2}}\right),\,\quad j=1,\ldots 7.$                                                                                                      The PrO posterior was run with $p=50$ particles for $10,000$ iterations. MALA was run for $4,000$ steps after $46,000$ warmup iterations, with $\mathsf{d}t=10^{-6}$. The MMD was estimated with $m=32$ Monte Carlo samples, and both models were fit with $\lambda_{n}=n^{1/2}$.
D.7 Time series cross-validation
Following Bürkner et al. (2020), suppose we have $T$ observations from a time series, $y_{i},\ldots,y_{T}$, and let $L$ be the minimum number of observations from the series that we will require before making predictions for future data. Then we want to evaluate the predictive performance of $M$ one-step-ahead predictions made by the model, starting with its performance on the $L+1$-th datum: $\int P_{\theta}(y_{L+1})\mathsf{d}Q_{n}(\theta\mid y_{1:L})$. For instance, the leave-future-out cross-validated one-step-ahead log score is given as $\sum_{i=L}^{T-1}\log\int P_{\theta}(y_{i+1}\mid y_{1:i})\mathsf{d}Q_{n}(\theta\mid y_{1:i}),$ and we compute the $\operatorname{MMD}^{2}$ and $\operatorname{RMSE}$ similarly: fitting the model on $i$ data and evaluating on the the $i+1$-th datum because incrementing the index $i$ by one until we have no more observations on which to evaluate. This procedure is elegantly visualised in Figure 1 of Bürkner et al. (2020).
D.8 Conditional auto-regression example
The PrO posterior under the MMD was run with $p=16$ particles for $2,000$ iterations. MALA was run for $4,000$ steps after $96,000$ warmup iterations, with $\mathsf{d}t=10^{-6}$. The MMD was estimated with $m=32$ Monte Carlo samples, and both models were fit with $\lambda_{n}=n^{1/2}$. In the case of the log score, the Gibbs posterior is fit with the same hyper-parameters as described above, while the PrO posterior is now fit with $p=128$ particles with $4,000$ iterations. Refer to caption Figure 15: The empirical data distribution of the median housing value by census tract.
Appendix E Computation
E.1 Examples: MMD and logarithmic score
Our above development recovers the algorithm of Shen et al. (2025) for the special case of squared MMD. Here, elementary calculations show that for $\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta;x)=\langle\mu(P_{\vartheta})-\mu(\updelta_{x}),\mu(P_{\theta})-\mu(\updelta_{x})\rangle_{\mathcal{H}}$ and $\nabla_{1}\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta;x)$ denoting its gradient with respect to $\vartheta$,                                                                                                     $\mathcal{W}(Q)[\vartheta]$ ${}={}$ $\frac{1}{n}\sum_{i=1}^{n}\int\nabla_{1}\mathsf{L}_{\operatorname{MMD}}(\vartheta,\theta;x_{i})\;\mathsf{d}Q(\theta),$                                                                                                      (34) so that the evolution equations take the form                                                                                                     $\mathsf{d}\vartheta_{t}^{(j)}=-\left\{\frac{1}{(p-1)}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{p}\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\nabla_{1}\mathsf{L}_{\operatorname{MMD}}(\vartheta_{t}^{(j)},\vartheta_{t}^{(\ell)};x_{i})-\nabla_{\vartheta}\log\mathsf{d}\Pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t}^{(j)}.$                                                                                                      (35) For the logarithmic scoring rule, one instead finds that                                                                                                     $\mathcal{W}(q)[\vartheta]$ ${}={}$ $\frac{1}{n}\sum_{i=1}^{n}\frac{\nabla_{\vartheta}\,\mathsf{d}P_{\vartheta}(x_{i})}{\int\mathsf{d}P_{\theta}(x_{i})\mathsf{d}Q(\theta)},$                                                                                                      (36) so that the evolution equations are recovered as                                                                                                     $\mathsf{d}\vartheta_{t}^{(j)}=-\left\{\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\frac{\nabla_{\vartheta}\,\mathsf{d}P_{\vartheta_{t}^{(j)}}(x_{i})}{\frac{1}{(p-1)}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{p}\mathsf{d}P_{\vartheta_{t}^{(\ell)}}(x_{i})}-\nabla_{\vartheta}\log\mathsf{d}\Pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t}^{(j)}.$                                                                                                      (37)
E.2 Practicalities and Implementation
To provide intuition about how the sampling algorithm described in this section differs from more conventional Bayesian computation procedures, Figure 12 compares it to the Metropolis adjusted Langevin algorithm (MALA; Roberts and Tweedie, 1996; Roberts and Rosenthal, 1998) on several normal location examples. While the $p=32$ particles which target the PrO posterior are initialised randomly from the prior, they soon repel each other, and lead to a multi-modal distribution. In contrast, the single particle evolved according to MALA bounces around a singular mode. While the interacting particle approach has the distinct advantage of being able to sample from any PrO posterior, it introduces several hyper-parameters that may require careful tuning in practice. The most important of these are the step sizes for the discretisation of the underlying SDE, the number of particles $p$, as well as the exact choice of $\lambda_{n}$. Here, the choice of step size and the number of particles $p$ constitute a direct trade-off between approximation quality and computational efficiency: choosing the step size too small or the number of particles too large will in principle increase accuracy, but do so at the cost of exploding computational cost and substantively slowed convergence. Determining what constituted an appropriate step size in particular, is non-trivial and an active area of research. Concurrently, Sharrock and Nemeth (2025) developed an adaptive step size schedule for sampling form mean-field Langevin dynamics by adaptively tuning the step size, which they call FUSE. We set the step size adaptively according to FUSE in all experiments. While the theoretical result those authors showed do not extend to the PrO posterior, our empirical results give promise that similar results are attainable. We leave this as an area of future research. The relationship between the dimensionality of $\Theta$ and the number of particles required for consistent estimation has been investigated in the related family of Stein variational gradient descent algorithms by Banerjee et al. (2025, Theorems 1 and 4). While those authors found that the number of particles must scale polynomially or exponentially with the dimension of the problem, we observed that when the step size is chosen with FUSE that much fewer particles are required. While little is known about choosing these hyper-parameters in our setting, recent work by Chazal et al. (2025) may finally provide a principled way for diagnosing this (see, e.g., Appendix D.4.1 for a numerical demonstration). In contrast, we have clear ideas on how $\lambda_{n}$ should be set from the theory developed in Section 3: for the case covered in Theorem 1 this is means setting $\lambda_{n}\asymp n^{1/2}/\log(n)$.